3) A 10 year zero coupon bond maturing at $1000 is callable at the 7 year point for $900.

a) What price should an investor pay if they want a (annual effective) yield of at least 5%?

b) What will their yield be if the bond is called early?
Answer #1

Part A:

Price of Bond = PV of CFs from it.

= CF * PVF (r%. n)

= $ 1000 * PVF(5%,10)

= $ 1000 * 0.6139

= $ 613.91

Part B:

Yield of Zero Coupon Bond = [ Maturity Value / Price of Bond ] ^ (1/n ) - 1

= [ 900 / 613.91 ] ^ ( 1 / 7 ) - 1

= 1.4660 ^ ( 1/7) - 1

= 1.0562 - 1

= 0.0562 i.e 5.62%